OpenStudy (freckles):

Evaluate \[\int_0^1(1-2x)(x-x^2)^{2014} dx\]

OpenStudy (freckles):

I was playing with @ganeshie8 's question and come across something weird.... I tried a substitution. u=x-x^2 but this would make both upper and lower limit 0 But for some reason I'm not convinced the integral should evaluate to 0.

OpenStudy (freckles):

I was trying to look for a non-beta/gamma way.

OpenStudy (freckles):

http://openstudy.com/study#/updates/54a59652e4b054f0c3b68124 This was @ganeshie8 's post by the way.

OpenStudy (turingtest):

The graph of \(f(x)=(1-2x)(x-x^2)^{2014}\) has 180 degree symmetry about x=1/2\[f(x+\frac12)=(1-2x-1)(x+\frac12-x^2-x-\frac14)=-2x(\frac14-x^2)\]In other words, it is "odd about x=1/2" (if I'm not just making up that terminology) hence any integral on an interval symmetric about x=1/2 will be zero.

OpenStudy (turingtest):

i forgot the exponent, but I doubt that causes confusion

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{0}^{1}~(1-2x)(x-x^2)^{2014}dx}\) derivative of u, is sitting inside there, nice:)

OpenStudy (solomonzelman):

I mean if I set: u=x-x^2

OpenStudy (solomonzelman):

I will let other people do this, but nice integral, I like this:)

ganeshie8 (ganeshie8):

wana see some action movie

OpenStudy (turingtest):

haha yeah software doesn't like huge exponents

ganeshie8 (ganeshie8):

another way to look at TT's reply :) : \[\int\limits_{0}^{1}(1-2x)(x-x^2)^{2014}dx \] substitute \(u = 1/2-x \implies 2u = 1-2x \) and \(x-x^2 = x(1-x) = 1/4-u^2\) the integral becomes \[\int\limits_{-1/2}^{1/2} (2u)(1/4-u^2)^{2014}\,du = 0\]

OpenStudy (turingtest):

that's what I wanted to write^, but I kept messing it up

ganeshie8 (ganeshie8):

guess we are just shifting the graph to left by 1/2 units so that it is symmetric about origin

OpenStudy (freckles):

ok yes I get it \[h(x)=x(x-x^2)^{2014}\] is "odd about zero" and \[f(x)=(1-2x)(x-x^2)^{2014}\] is just a translation of the graph h and it should be "odd about 1/2" and nice picture

OpenStudy (solomonzelman):

don't get why can't we just set u=x-x^2, lol. it is made for that.

OpenStudy (solomonzelman):

it's pretty basic:)

OpenStudy (solomonzelman):

someone is trying to increase engagement score? jk

OpenStudy (turingtest):

freckels had the answer, but realized that the substitution u-x-x^2 make the bounds both zero, and they wanted intuition as to why the integral should be zero.

OpenStudy (solomonzelman):

sure

OpenStudy (turingtest):

u=x-x^2 *

OpenStudy (solomonzelman):

I realize the typo:D